OpenStudy (anonymous):

6b^2-5b=4 solve the equation and check your solution

OpenStudy (anonymous):

To solve it, you must get everything to one side so the equation equals zero. 6b^2 - 5b - 4 = 0 Then factor it.

OpenStudy (anonymous):

Do you know how to factor that?

OpenStudy (anonymous):

no thats as far as ive gotten

OpenStudy (anonymous):

is the 6b^2-5b-4=0

OpenStudy (anonymous):

Do you know if you use trial and error, bottoms up, or what method you usually use?

OpenStudy (anonymous):

foil

OpenStudy (anonymous):

Foil would be to check it

OpenStudy (anonymous):

Factoring is taking it apart. We'll use trial and error. Multiply the first and last numbers and you get -24

OpenStudy (anonymous):

What multiplies to be -24 and adds to be -5 (the middle term)

OpenStudy (anonymous):

3 and 8

OpenStudy (anonymous):

Yeah, 3 and -8. So you know your "inner" and outter terms have to equal 3 and

OpenStudy (anonymous):

6 factors to be 2 * 3 or 1 * 6 and 4 can be 2 * 2 or 4 *1

OpenStudy (anonymous):

You know you have to have "3" for one of your numbers, so you know it's set up like this.... (3x + ____) (2x + ____)

OpenStudy (anonymous):

Since you know you need a 3. that last blank must be a positive 1. (3x + ___)(2x + 1) . That makes your outter terms equal 3.

OpenStudy (anonymous):

Therefore your inner terms must equal -8. So what would need to go in that blank?

OpenStudy (anonymous):

-4

OpenStudy (anonymous):

Yup. So that gives you (3x - 4)(2x + 1) Then to check it, just FOIL it back together.

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

you're welcome

OpenStudy (anonymous):

wait shouldnt the answer come out to be x= and then like 2 fractions cause on the last one that i did that asked for the equation to be solved and check solution it came out to be n=-1/3 and 2/5

OpenStudy (anonymous):

Yes, just set each ( ) equal to zero and solve. The answers should be x = 4/3 and x = -1/2